OpenStudy (anonymous):

Find the inverse of \[h(x)=8sqrt{8x-3}-1\] where \[x ge \frac{3}{8}\]

OpenStudy (anonymous):

\[h(x)=8 \sqrt{8x-3}-1\] where \[x \ge \frac{3}{8}\] and @saifoo.khan I know what you're thinking... I post a question just like this... but I just want to make sure am doing the right thing... :D.

OpenStudy (anonymous):

And... I have not been given one with certain conditions in class...

OpenStudy (saifoo.khan):

I'm not sure i can solve this one. sorry.

OpenStudy (saifoo.khan):

But i will try.

OpenStudy (saifoo.khan):

(y+1)/8 = sqrt(8x-3) take square on both sides.. \[(\frac{y+1}{8})^2 = 8x-3\]

OpenStudy (saifoo.khan):

@dpaInc , HELP!

OpenStudy (anonymous):

I got... \[f^{-1}(x)=\frac{1}{8}(\frac{x+1}{2})^2+\frac{3}{8}\] but I don't know if I have to clean it up...

OpenStudy (saifoo.khan):

it's 8 sqrt or 2 sqrt ?

OpenStudy (anonymous):

8

OpenStudy (saifoo.khan):

Then how you got (x+1)/2 ?

OpenStudy (anonymous):

Oh that is suppose to be an 8... Idk why am making up numbers...

OpenStudy (saifoo.khan):

Then following up, \[\frac{(y+1)^2}{64} +3 = 8x\]

OpenStudy (saifoo.khan):

brb 10mins

OpenStudy (anonymous):

Seems as if you didn't exchange x and y...

OpenStudy (apoorvk):

Hmm, where are we stuck now?

OpenStudy (anonymous):

I got: \[\frac{y^{2} + 2x + 193}{512}\] however, I'm not sure what the purpose of the x ≥ 3/8 part is about...

OpenStudy (apoorvk):

well yeah, x \(\ge\) 3/8 since restrictions apply for the original equation for the part under the root. otherwise you're right @Calcmathlete , just need to interchange 'y' and 'x' now.

OpenStudy (anonymous):

Oh alright. \[f(x)^{-1} = \frac{x^{2} + 2x + 193}{512}\]

OpenStudy (apoorvk):

Almost! just the wee tiny winy thing! :p \(\large f^{-1}(x) = \frac{x^{2} + 2x + 193}{512}\)

OpenStudy (anonymous):

lol. Slight mix up :)

OpenStudy (apoorvk):

no worries! :D

OpenStudy (anonymous):

Oh thanks guys! I got the same thing :)

OpenStudy (anonymous):

np :)

OpenStudy (apoorvk):

Yo!